OpenStudy (anonymous):

A car dealer will sell you a used car for $7,635 with $635 down and payments of $267.36 per month for 36 months. What is the simple interest rate?

jimthompson5910 (jim_thompson5910):

The dealer will finance (ie loan you) 7635 - 635 = 7000 dollars So P = 7000 You pay 267.67 per month for 36 months. So you pay 267.67*36 = 9636.12 over the total life of the loan (excluding down payment) This means A = 9636.12 The rate is unknown. So we'll call it x. So r = x The time (in years) is 36/12 = 3. So t = 3

jimthompson5910 (jim_thompson5910):

Plug all this info into A = P(1+rt) and solve for x

OpenStudy (anonymous):

Wouldn't you just take 9636.13-7000 to find the interest?

jimthompson5910 (jim_thompson5910):

that's another way you can do it

jimthompson5910 (jim_thompson5910):

9636.12-7000 = 2636.12 is the amount of interest you pay

OpenStudy (anonymous):

Is that different from simple interest?

jimthompson5910 (jim_thompson5910):

no it's just taking another route to the same answer

OpenStudy (anonymous):

That' what I thought. Thanks

jimthompson5910 (jim_thompson5910):

you're welcome